How many grams of aluminum chloride will be produced from 255 moles of Cl2


1
Expert's answer
2021-03-23T07:25:15-0400

Solution:

The balanced chemical equation:

2Al + 3Cl2 → 2AlCl3

According to the equation above: n(Cl2)/3 = n(AlCl3)/2


n(AlCl3) = 2 × n(Cl2) / 3 = (2 × 255 mol) / 3 = 170 mol


Moles of AlCl3 = 170 mol

Mass of AlCl3 = Moles of AlCl3 × Molar mass of AlCl3

The molar mass of AlCl3 is 133.5 g mol-1.

Hence,

Mass of AlCl3 = 170 mol × 133.5 g mol-1 = 22695 g

Mass of AlCl3 = 22695 g


OR (short form solution):

(255 mol Cl2) × (2 mol AlCl3 / 3 mol Cl2) × (133.5 g AlCl3 / 1 mol AlCl3) = 22695 g AlCl3


Answer: 22695 grams of aluminum chloride (AlCl3) will be produced.